OpenStudy (anonymous):

Given sin (theta) = 21/29... find sin cos and tan (2 theta)

OpenStudy (loser66):

\(sin(\theta)=\dfrac{opposite}{hypotenuse}\) how about cos(theta)?

OpenStudy (anonymous):

adj / hyp

OpenStudy (loser66):

ok, let them there. Now, pythagorean  Can you put numbers on the triangle?

OpenStudy (loser66):

on that triangle, which side is hypotenuse?

OpenStudy (anonymous):

left side is hyp bottom is adj and right side is opp

OpenStudy (loser66):

you use draw box below, click on it, then a box will pop up. On the right corner, you will see "copy previous drawing", click on it, then scroll up to see my picture, click it to get the pic, then put the numbers in, then post

OpenStudy (loser66):



OpenStudy (loser66):

hence, the adj side can be calculated by Pythagorean. Do you know him?

OpenStudy (loser66):

don't get??

OpenStudy (anonymous):

oh a2 + b2 = c2

OpenStudy (loser66):

Mr. Pythagore said : In this case, \(adj = \sqrt{29^2-21^2}=??\)